OpenStudy (anonymous):

y varies directly as the square of x. When x = 5, y = 5. Find y when x = 10. (4 points) Answer y = 10 y = 20 y = 50 y = 100

OpenStudy (anonymous):

no medal for my last one?

OpenStudy (anonymous):

i gave it to you .

OpenStudy (anonymous):

well the answer is y=10

OpenStudy (anonymous):

oh wait no, sorry

OpenStudy (anonymous):

i just realized it said the square

OpenStudy (anonymous):

so thats not the answer ?

OpenStudy (anonymous):

no it's 20

OpenStudy (anonymous):

y = a x^2, with x=5 and y=5 that means a = 1/5

OpenStudy (anonymous):

imma give you a medal for every problem you do for me . (:

OpenStudy (anonymous):

y = 1/5 x^2 with x=10 gives you 20

OpenStudy (anonymous):

20